A solution of ethanol, CH3CH2OH, in water has a concentration of 1.40 m. Calculate the mass percent of ethanol.
Answer #1

The concentration = 1.40 m (molality)

that means 1.4 moles of ethanol is added in  1 Kg of water.

Thus mass of ethanol in solution = moles x molar mass = 1.40 mol x 46g/mol =64.4 g

mass of water = 1 kg =1000 g

mass of solution = mass of ethanol + mass of water = 64.4+ 1000 = 1064.4g

Therefore

% by mass of ethanol = mass of ethanol x 100/ mass of solution

= 64.4g x100/1064.4 g = 6.05 %